# Task-Oriented Automatic Fact-Checking with Frame-Semantics

### v2

@@ 4.2 OECD Statistics
The OECD provides a wealth of trustworthy, observational data on a diverse set of statistics which serve as a strong focal point to explore a wider range of semantic frames. For the OECD case study, we collected all of the data tables available on the OECD Data Explorer.  [^7] https://data-explorer.oecd.org/ We constructed an SQL database consisting of each data source, resulting in 434 tables with an average of 596,552 rows per table, totaling over 4.1 billion cells.
@@
To identify the relevant table(s) to fact-check a given claim, we use a RoBERTa-based semantic similarity model Liu et al. (2019). To represent the table, we encode the table’s name and description as text. To minimize the likelihood of missing relevant tables, we retrieve the five tables most similar to the extracted frame element from the claim.
@@
Querying the retrieved tables requires understanding of the tables, columns, and the values within them. To do this, we encode each table’s name, all of its columns, and several representative sample values from each column into text. We then utilize GPT-4o to understand the requirements for verifying a claim and generate Python code to query the retrieved table accordingly. We choose this approach, as opposed to standard text-to-SQL approaches because it allows for the execution of multiple simple queries using programmed logic, rather than relying on text-to-SQL generation which can result in very complex queries for simple claims.
We represent the table schema as SQL code, following best practices from previous work Gao et al. (2024). Alongside the schema, we include representative example values from the database for each column. Our implementation is similar to Nahid and Rafiei (2024); however, we address critical limitations to support the high-volume OECD data. Encoded columns with fewer than five unique values include all values, while columns with more than 100 unique values are randomly sampled for ten representative values. For other columns we use the same RoBERTa-based semantic similarity model to select the top ten most relevant values based on the claim. These queries often result in a large number of cells, so we instruct GPT-4o to refine the query by filtering out irrelevant columns, focusing on those most critical for fact-checking the claim. The prompts for this process can be found in Appendix E.3.
@@ 5.1 Datasets
Arslan et al. (2020) introduced 11 manually defined semantic frames to extend the long-running Berkeley FrameNet project Baker et al. (1998). Annotations for 936 sentences containing 1,029 frame-evoking targets and 3,570 frame elements are included in their dataset along with the newly-defined frames. These frames can be found in Table 6 in the Appendix.
@@
For our analysis, we utilized a dataset of 21,024 PolitiFact fact-check articles (i.e., fact-checks) collected as of April 2022. Each article contains a claim, a detailed fact-checking analysis, a verdict, and a list of sources. To focus on voting-related claims, we extracted 1,552 (7.4%) fact-checks that mention some form of “vote.” From this subset, we manually identified 79 fact-checks that cite congress.gov in the sources and evoke the Vote frame. For each voting claim, we manually verified the bills referenced within the fact-check are related to the claim to ensure the accuracy of our dataset. Similarly, we collected 68 fact-checks that cite oecd.org for our analysis of OECD-related claims. Each OECD claim was manually verified and mapped to the database table which can be used to fact-check it. Additional details are provided in Appendix A.
@@
Table 1: Performance of frame-semantic parsing model on fact-checking frames alongside simple baseline.
@@
Table 2: Distribution of semantic frames in PolitiFact.
@@ 5.2 Frame-Semantic Parsing
For frame-semantic parsing, we combined the frame identification model developed by Devasier et al. (2024) with a frame element identification model based on AGED Zheng et al. (2023). By combining these state-of-the-art approaches, we unified frame and frame element identification into a single RoBERTa-based model (additional details in Appendix C.4). We fine-tuned two of these frame-semantic parsing models, one for each case study, using their respective frames from the training data in Section 5.1. The model used by our system is then selected based on the candidate targets in the sentence (extracted using the approach in Devasier et al., 2024). To evaluate the performance of our model for the case studies, we compared it with a baseline GPT-4o-mini model using OpenAI’s structured generation. The prompt used for this model can be found in Appendix E.1. We evaluated each model using exact match accuracy on frame identification and argument identification. These results are presented in Table 1.
@@ 5.3 PolitiFact Survey
To assess the broader applicability of our frame-semantics-based fact-checking system, we conducted a comprehensive coverage analysis. Because the annotations available for the fact-checking frames are limited, we employed a zero-shot GPT-4o-mini model (detailed in Appendix E.1) to identify frames evoked in the PolitiFact fact-check corpus. Claims that did not trigger any of our studied frames were classified as “None.” The distribution of frames is presented in Table 2.
Our findings show that GPT-4o-mini may predict more frames than there actually are. This is because the annotated dataset evokes an average of 1.1 frames per claim while GPT-4o-mini predicted an average of 2.1 frames per claim. We believe this is because GPT-4o-mini predicts frames for lexical units–words which evoke frames–that are not defined by Arslan et al. (2020).color=yellow]This last sentence appears to suggest GPT can find frames there were missed in Fatma’s dataset. But why did you say it “predict more than there actually are”? — This isn’t really known 100% the cause could be that the model over-predicts in a bad way or it could be because the model recognizes frames for more LUs than are defined. It seems to be the latter based on what ive seen.
@@ 5.4 Evidence Retrieval
To retrieve relevant information for fact-checking, we first evaluated different semantic similarity models. Table 3 presents the embedding models we compared, including Stella Zhang et al. (2025), RoBERTa Liu et al. (2019), DistilBERT-TAS-B Hofstätter et al. (2021), OpenAI’s text-embedding-3, 999 https://platform.openai.com/docs/models and BM25. 101010 https://github.com/dorianbrown/rank_bm25 color=yellow]In Table 4, font size of different models are different; some R@K is italic, others not. Table 1 and Table 3/4 use different font styles.
@@
We explored two different methods for constructing retrieval queries: using the complete claim text and using only the frame elements (FEs) extracted based on the evoked frame (detailed in Appendix B). We evaluated these query representations using the best-performing similarity models identified in Table 3.
@@ 5.5 Fact Verification
We evaluated our fact verification system by comparing its predicted verdicts against the ground truth verdicts from PolitiFact fact-checks (Table 5). For consistency in our analysis, we consolidated PolitiFact’s "pants on fire" rating with "false".
@@ 6 Conclusion
Another limitation is that our system’s effectiveness is constrained by the availability of reliable, manually curated data. For claims that lack relevant data in our databases (e.g., OECD tables or U.S. congressional records), the system cannot retrieve evidence, which limits its fact-checking capabilities. This limitation demonstrates the importance of expanding data sources and improving coverage in future iterations of the system.
@@
Lastly, the over-dependence on specific sources like PolitiFact or congress.gov could limit the tool’s scope. Ethical development of such systems requires attention to these risks to ensure fairness, accuracy, and social responsibility in their use.
@@ Appendix A Data Quality
Model Parameters URL
msmarco-distilbert-base-tas-b 66M msmarco-distilbert-base-tas-b
msmarco-roberta-base-v2 125M msmarco-roberta-base-v2
msmarco-roberta-base-v3 125M msmarco-roberta-base-v3
stella_en_1.5B_v5 1.5B stella_en_1.5B_v5
text-embedding-3-large unknown text-embedding-3-large
Table 7: Models and their corresponding Hugging Face URLs.
@@ C.1 Computational Resources
We trained our frame-semantic parser on an NVIDIA GTX 1080 Ti GPU. A single training run took approximately 6 hours. No hyperparameter tuning was performed, and the model was trained using default settings unless explicitly stated. The OECD database requires roughly 85GB of storage and the voting records requires 1GB of storage. color=cyan]we don’t train any other models.
@@ C.4 Models
Our frame-semantic parsing model is an adaptation of AGED, augmented with the frame classification layer and strategy from Devasier et al. (2024). We fine-tuned this model using a joint loss function to improve frame prediction accuracy. Instead of BERT, as originally used in AGED, we employ RoBERTa for its improved contextual embeddings, which have been shown to enhance performance on frame-based semantic parsing tasks.
@@
Our agent matching function retrieves the BioGuide ID of a political figure from a database. It extracts the agent’s name using frame element spans (FEs) and checks if it is a known presidential nickname, replacing it with the full name if applicable. If not, it uses a SpaCy NER model to extract PERSON named entities. The function first attempts an exact match in the members table, then a fuzzy match using SQL’s LIKE operator. If neither succeeds, it constructs a multi-condition query, incorporating nickname expansions where applicable.
@@ D.1 Vote-Claim Alignment
Table 8 presents two sample claims from the voting dataset along with their top three retrieved bills. For each retrieved bill, we include the computed vote-claim alignment score, which indicates the relevance of the bill to the claim. This alignment process plays a crucial role in determining the factual accuracy of voting-related claims.color=red]I don’t any score in the table.
@@ D.2 Fact Verification Predictions
For voting-related claims, we provide two examples where our model predicts the veracity of claims based on retrieved bills. Table 8 includes the claims, retrieved evidence, and the final prediction score.color=red]no score

### v3

@@ 4.2 OECD Statistics
The OECD provides a wealth of trustworthy, observational data on a diverse set of statistics which serve as a strong focal point to explore a wider range of semantic frames. For example, there are statistics related to health (including healthcare coverage, health risk factors, pharmaceutical markets, etc.), financial information (including pension assets, stocks and investments, employment earnings, unemployment rates, etc.), scientific data (air pollution, greenhouse gas emissions, exposure to droughts/floods/wildfires, energy consumption, plastic leakage, etc.), and many more topics. With this dataset and the wider range of frames, we drastically expand both the scope of topics (leveraging the OECD data) and the types of claims (leveraging the frames) our system can check. For the OECD case study, we collected all of the data tables available on the OECD Data Explorer. [^7] https://data-explorer.oecd.org/ We constructed an SQL database consisting of each data source, resulting in 434 tables with an average of 596,552 rows per table, totaling over 4.1 billion cells.
To evaluate our system, like the previous case study, we constructed an evaluation dataset by extracting PolitiFact fact-checks before April 2022 that cite oecd.org in their fact-check sources. After manual verification, we collected 68 fact-checks, along with their corresponding OECD data table, to form the evaluation set.
@@
To identify the relevant table(s) to fact-check a given claim, we use a RoBERTa-based semantic similarity model Liu et al. (2019). To represent the table, we encode the table’s name and description as text. To minimize the likelihood of missing relevant tables, we retrieve the five tables most similar to the extracted frame element (shown for each frame in Table 6) from the claim.
@@
Querying the retrieved tables requires understanding of the tables, columns, and the values within them. To do this, we encode each table’s name, all of its columns, and several representative sample values from each column into text. We then utilize GPT-4o to understand the requirements for verifying a claim and generate Python code to query the retrieved table(s) accordingly. We choose this approach, as opposed to standard text-to-SQL approaches Lei et al. (2024) because it allows for the execution of multiple simple queries using programmed logic, rather than relying on text-to-SQL generation which can result in very complex queries for simple claims.
We represent the table schema as SQL code, following best practices from previous work Gao et al. (2024). Alongside the schema, we include representative example values from the database for each column. Our implementation is similar to Nahid and Rafiei (2024); however, we address critical limitations to support the high-volume OECD data. Encoded columns with fewer than five distinct values include all values, while columns with more than 100 distinct values are randomly sampled for ten representative values. For other columns we use the same RoBERTa-based semantic similarity model from the relevant table identification to select the top ten most relevant values based on the claim. Querying the databases often result in a large number of cells, so we instruct GPT-4o to refine the query by filtering out irrelevant columns, focusing on those most critical for fact-checking the claim. The prompts for this process can be found in Appendix E.3.
@@ 5.1 Datasets
Arslan et al. (2020) introduced 11 manually defined semantic frames to extend the long-running Berkeley FrameNet project Baker et al. (1998). Annotations for 936 sentences containing 1,029 frame-evoking targets and 3,570 frame elements are included in their dataset along with the newly-defined frames. These frames, along with the frame elements used for identifying relevant tables, can be found in Table 6 in the Appendix.
@@
For our analysis, we utilized a dataset of 21,024 PolitiFact fact-check articles (i.e., fact-checks) collected as of April 2022. Each article contains a claim, a detailed fact-checking analysis, a verdict, and a list of sources used in the fact-check. To focus on voting-related claims, we extracted 1,552 (7.4%) fact-checks that mention some form of “vote.” From this subset, we manually identified 79 fact-checks that cite congress.gov in the sources and evoke the Vote frame. For each voting claim, we manually verified the bills referenced within the fact-check are related to the claim to ensure the accuracy of our dataset. Similarly, we collected 68 fact-checks that cite oecd.org for our analysis of OECD-related claims. Each OECD claim was manually verified and mapped to the database table which can be used to fact-check it. Additional details are provided in Appendix A.
@@
Table 1: Performance of frame-semantic parsing model on fact-checking frames compared with LLM and random choice baselines.
@@
Table 2: Distribution of semantic frame predictions in fact-checking articles using GPT-4o-mini.
@@ 5.2 Frame-Semantic Parsing
For frame-semantic parsing, we combined the frame identification model developed by Devasier et al. (2024) with a frame element identification model based on AGED Zheng et al. (2023). By combining these state-of-the-art approaches, we unified frame and frame element identification into a single RoBERTa-based model (additional details in Appendix C.4). We fine-tuned two of these frame-semantic parsing models, one for each case study, using their respective frames from the training data in Section 5.1. The model used by our system is then selected based on the candidate targets in the sentence (extracted using the approach in Devasier et al., 2024). To evaluate the performance of our model for the case studies, we compared it with a baseline GPT-4o-mini model using OpenAI’s structured generation. The prompt used for this model can be found in Appendix E.1. We evaluated each model using exact match accuracy on frame identification and argument identification (predicting the frame elements of the evoked frame). These results are presented in Table 1.
@@ 5.3 PolitiFact Survey
To assess the broader applicability of our frame-semantics-based fact-checking system, we conducted a comprehensive coverage analysis. For this analysis, we used the frames discussed in Arslan et al. (2020), including the 11 newly defined frames from their work. Because the annotations available for the fact-checking frames are limited, we employed a zero-shot GPT-4o-mini model (detailed in Appendix E.1) to identify frames evoked in the PolitiFact fact-check corpus. Claims that did not trigger any of our studied frames were classified as “None.” The distribution of frames is presented in Table 2.
Applying GPT-4o-mini resulted in 45,772 predicted frames, indicating that it may over-predict relative to the annotated data. While the gold dataset averages 1.1 frames per claim, GPT-4o-mini predicts 2.1 frames on average. This likely stems from GPT-4o-mini assigning frames to lexical units—words that can evoke frames—that were not defined by Arslan et al. (2020).
@@
distilroberta-base Vote 0.114
all-mpnet-base-v2 Vote 0.115
@@
multi-qa-mpnet-base Vote 0.143
stella_en_400M_v5 Vote 0.144
@@ 5.4 Evidence Retrieval
To retrieve relevant information for fact-checking, we first evaluated different semantic similarity models. We experimented with several models from the MTEB leaderboard.  [^9] https://huggingface.co/spaces/mteb/leaderboard Table 3 presents the best-performing models, including Stella Zhang et al. (2025), RoBERTa Liu et al. (2019), DistilBERT-TAS-B Hofstätter et al. (2021), OpenAI’s text-embedding-3, 101010 https://platform.openai.com/docs/models and BM25. 111111 https://github.com/dorianbrown/rank_bm25
@@
We explored two different methods for constructing retrieval queries: using the complete claim text and using only the frame elements (FEs) extracted based on the evoked frame (detailed in Appendix B). For each frame, we selected specific frame elements (Table 6) which are most relevant for evidence retrieval. We evaluated these query representations using the best-performing similarity models identified in Table 3.
@@ 5.5 Fact Verification
We evaluated our fact verification system by comparing its predicted verdicts against the ground truth verdicts from PolitiFact fact-checks (Table 5). For consistency in our analysis, we consolidated PolitiFact’s “pants on fire” rating with “false”.
@@ 6 Conclusion
Another limitation is that our system’s effectiveness is constrained by the availability of reliable, manually curated data. For claims that lack relevant data in our databases (i.e., OECD tables or U.S. congressional records), the system cannot retrieve evidence, which limits its fact-checking capabilities. This limitation demonstrates the importance of expanding data sources and improving coverage in future iterations of the system.
Finally, the method relies on existing FrameNet frames. FrameNet versions for languages other than English are often incomplete or entirely unavailable. As a result, the current method may not be effective for fact-checking in non-English languages.
@@
Lastly, the over-dependence on specific sources such as PolitiFact or congress.gov could limit the tool’s scope. Ethical development of such systems requires attention to these risks to ensure fairness, accuracy, and social responsibility in their use.
Acknowledgments
This work is partially supported by the National Science Foundation under Grant 2346261. We also extend our gratitude to the Texas Advanced Computing Center (TACC) for providing compute resources used in this work’s experimentation.
@@ Appendix A Data Quality
Model Parameters
msmarco-distilbert-base-tas-b 66M
msmarco-roberta-base-v2 125M
msmarco-roberta-base-v3 125M
stella_en_1.5B_v5 1.5B
text-embedding-3-large unknown
Table 7: Models and their parameter sizes.
@@ C.1 Computational Resources
We trained our frame-semantic parser on an NVIDIA GTX 1080 Ti GPU. A single training run took approximately 6 hours. No hyperparameter tuning was performed, and the model was trained using default settings unless explicitly stated. The OECD database requires roughly 85GB of storage and the voting records requires 1GB of storage.
@@ C.4 Models
Our frame-semantic parsing model builds on AGED, which achieves state-of-the-art performance in identifying the spans of frame elements for a given frame in a sentence. While AGED focuses on span detection, Devasier et al. (2024) propose a complementary training strategy that improves frame identification by incorporating additional negative examples during training. Both models use similar underlying architectures, which allowed us to integrate their strengths. Specifically, we augmented AGED with an additional classification layer to predict the correct frame and adopted the robust training procedure introduced by Devasier et al. For our base encoder, we used RoBERTa instead of BERT, as it provides stronger contextual representations. The resulting model demonstrated strong empirical performance, so we retained this architecture for all downstream tasks.
@@
Our agent matching function retrieves the BioGuide ID  [^13] https://www.congress.gov/help/field-values/member-bioguide-ids of a political figure from the voting records database. It extracts the agent’s name using frame element spans and checks if it is a known political nickname, replacing it with the full name if applicable. If not, it uses a SpaCy NER model [^14] https://spacy.io/api/entityrecognizer to extract PERSON named entities. The function first attempts an exact match then a fuzzy match using SQL’s LIKE operator if the exact match fails. If neither succeeds, it constructs a multi-condition query, incorporating nickname expansions where applicable.
@@ D.1 Vote-Claim Alignment
Table 8 presents two sample claims from the voting dataset along with their top three retrieved bills. For each retrieved bill, we include the computed vote-claim alignment classification (one of Supports, Refutes, Irrelevant). This alignment process plays a crucial role in determining the factual accuracy of voting-related claims.
@@ D.2 Fact Verification Predictions
For voting-related claims, we provide two examples where our model predicts the veracity of claims based on retrieved bills. Table 8 includes the claims, retrieved evidence, and the final prediction class.